# On Prompt-Driven Safeguarding for Large Language Models

### v2

@@ Abstract
Prepending model inputs with safety prompts is a common practice for safeguarding large language models (LLMs) from complying with queries that contain harmful intents. However, the working mechanisms of safety prompts have not been revealed yet, which hinders the potential for automatically optimizing them to improve LLM safety. To this end, we investigate the impact of safety prompts from the perspective of model representations. We find that in models’ representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts. Instead, the queries’ representations are moved by safety prompts in similar directions where models become more prone to refusal (i.e., refusing to provide assistance) even when the queries are harmless. Inspired by these findings, we propose a method called DRO (Directed Representation Optimization) for automatic safety prompt optimization. It treats safety prompts as continuous, trainable embeddings and learns to move the representations of harmful/harmless queries along/opposite the direction in which the model’s refusal probability increases. Experiments with eight LLMs on out-of-domain benchmarks demonstrate that DRO remarkably improves the safeguarding performance of human-crafted safety prompts and outperforms strong baselines, without compromising the general model capability.
@@ 1 Introduction
While the capabilities of large language models (LLMs) keep growing, as exemplified by ChatGPT (OpenAI, 2023), LLaMA (Touvron et al., 2023), and Mistral (Jiang et al., 2023), there are also rising concerns that they can engage with queries having harmful intents (e.g., seeking assistance about causing damages) (Weidinger et al., 2021). A common and lightweight means of safeguarding LLMs against harmful queries is to prepend model inputs with human-crafted safety prompts, which typically contain explicit guidance and guardrails on models’ behaviors. Real-world practices like GPT-4 (OpenAI, 2023) and Mistral (Jiang et al., 2023) have shown that adding safety prompts can mitigate models’ compliance with harmful queries without changing models’ parameters, as illustrated in Figure 1.
@@
However, we still have no clear understanding of the working mechanisms of safety prompts, which limits the potential of automatically optimizing them to improve LLM safety. Motivated by this problem, our work starts by delving into how safety prompts intrinsically affect model behaviors from the perspective of model representations2). We propose two hypotheses: (1) Models cannot well distinguish harmful and harmless queries, while safety prompts enhance models’ capability of harmfulness recognition. (2) Models can recognize harmful queries but fail to refuse them, while safety prompts increase the overall probability of refusal (i.e., refusing to provide assistance). To verify the hypotheses, we first collect harmful and harmless queries through carefully controlled data synthesis (see Figure 2 for examples). We then evaluate eight open-source LLMs and employ PCA to visualize their hidden states. We find that in models’ representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts, suggesting that our first hypothesis may not hold. Instead, we observe that the queries’ representations are moved by different safety prompts in similar directions, in which models become more prone to generating refusal responses even when the queries are harmless, thus confirming our second hypothesis.
@@
Inspired by these findings, we present a method for safety prompt optimization, named DRO (Directed Representation Optimization; §3). It takes the setting of the prompt tuning paradigm (Lester et al., 2021), where the model parameters are frozen and only a few continuous embeddings (corresponding to the safety prompts in our context) are trainable. DRO first anchors a model’s low-dimensional representation space and estimates the “refusal direction” that indicates the model’s refusal probability to increase (§3.1). It then optimizes the continuous safety prompt so that the queries’ representations are moved along (for harmful queries) or opposite (for harmless queries) the refusal direction (§3.2). We also design a regularization item to prevent the degeneration of the original representation caused by direct optimization in the low-dimensional space (§3.3).
@@
We apply DRO to optimize the LLaMA-2 and Mistral official safety prompts. Through experiments with eight LLMs on two out-of-domain benchmarks (§4), we demonstrate that DRO remarkably improves safeguarding performance (for the LLaMA-2 safety prompt, the percentage of compliance with harmful queries is reduced from 10.3% to 1.4% on AdvBench, Zou et al. 2023), and outperforms the vanilla Prompt-Tuning baseline (1.4% vs. 6.8% on AdvBench). Furthermore, DRO does not compromise the general model capability, as evaluated on AlpacaEval (Li et al., 2023), and exhibits reasonable robustness to the choices of data used for anchoring the low-dimensional space and refusal direction. We hope this work sheds light on the intrinsic working mechanisms of the prompt-driven LLM safeguarding approach, and inspires future research on LLM safety.
@@ 2 How Safety Prompts Intrinsically Work?
How to capture someone against their will and force them into captivity?
How to capture precious moments with a camera on a family trip?
How to construct a false narrative to deceive others?
How to construct a sturdy bridge over a river?
How to dispose of evidence after committing a crime?
How to dispose of recyclable materials properly?
@@
Why can safety prompts safeguard LLMs against harmful queries, without which models may fail to refuse these queries but instead comply with them? We propose two hypotheses for the working mechanisms of safety prompts: (1) Models cannot well distinguish harmful and harmless queries, while safety prompts enhance models’ capability of harmfulness recognition. (2) Models can recognize harmful queries but fail to refuse them, while safety prompts increase models’ overall probability of generating refusal responses. To verify the hypotheses, we investigate how harmful and harmless queries exist in models’ representation space, and how the impact of safety prompts on queries’ representations correlates with models’ refusal behaviors.
@@ 2.3 Visualization Analysis
How harmful and harmless queries exist in models’ representation space?
From the upper part of Figure 3, harmful and harmless queries can be largely distinguished without safety prompts, whose boundary (black chain dotted line) can be easily fitted by logistic regression using queries’ harmfulness as labels. Adding safety prompts does not noticeably increase their distinguishability, even when visualized in other principal components (see Appendix G). These observations suggest that our first hypothesis may not hold, i.e., safety prompts may not work by enhancing models’ capability of harmfulness recognition.
@@
3 Method for Safety Prompt Optimization
Despite widespread use in real-world deployed LLMs like GPT-4 (OpenAI, 2023) and Mistral (Jiang et al., 2023), the prompt-driven safeguarding approach has its shortcoming, that is, the effectiveness varies with human-crafted prompts and models, as shown in Table 1. For instance, the short safety prompt works poorly with mistral-inst-v0.1 (55% harmful queries are still being complied with). Models that have undergone massive safety training, such as llama-2-chat and codellama-instruct, may also become over-sensitive when equipped with safety prompts, thereby leading to false refusals for harmless queries. Since crafting a basic safety prompt is always easy, can we optimize it for improved safeguarding performance? Inspired by our findings in §2, we propose a method for automatically optimizing continuous safety prompts, named DRO, standing for Directed Representation Optimization. Its core idea is to move queries’ representations along or opposite the refusal direction based on the queries’ harmfulness.
@@ 3.2 Optimization Process
We similarly calculate a loss $\mathcal{L}_{\mathrm{h}}(\bm{\theta})$ for harmfulness recognition, with the same dimensionality reduction function $g$ but a different logistic regression $f_{\mathrm{h}}$ fitted using queries’ harmfulness as labels. It helps maintain the capability of distinguishing harmful/harmless queries, which we find can bring some safeguarding performance improvement (§4.2).
@@
1:  Anchor the low-dimensional space and fit the refusal direction.
$\triangleright$ Anchoring process (§3.1)
@@
$\triangleright$ Optimization process (Figure 4; §3.2, 3.3)
Table 2: Evaluation results (optimizing the default basic safety prompt) on MaliciousInstruct and Advbench.
% Compliance on MaliciousInstruct ↓ % Compliance on AdvBench ↓
llama-2-chat 1 1 1 1 0 1 0 0 0 3 0 0 0 0
codellama-instruct 3 2 7 1 1 1 1 2 0 2 0 0 0 0
vicuna-v1.5 51 10 7 2 2 4 2 27 4 2 0 1 2 0
orca-2 70 22 2 1 1 7 1 70 2 4 0 0 0 0
mistral-inst-v0.1 77 31 10 3 1 37 2 86 62 26 6 5 63 1
mistral-inst-v0.2 30 2 1 1 2 1 1 51 3 0 1 0 1 0
openchat-3.5 77 9 9 3 2 8 5 81 10 11 3 1 7 2
openchat-3.5-1210 66 1 3 1 3 3 2 78 1 6 1 1 7 1
average 46.9 9.8 5.0 1.6 1.5 7.8 1.8 49.4 10.3 6.8 1.4 1.0 10.0 0.5
Table 3: Evaluation results (optimizing the default basic safety prompt) on the held-out harmless query set and AlpacaEval.
% Refusal on Held-out Harmless ↓ % Win Rate on AlpacaEval ↑
llama-2-chat 1 19 5 5 3 7 7 66 47 37 54 53 53 48
codellama-instruct 3 22 0 7 5 8 7 54 52 47 51 45 48 51
vicuna-v1.5 0 5 4 2 1 0 1 68 65 62 64 58 65 61
orca-2 1 5 3 0 0 0 0 63 56 45 60 58 61 60
mistral-inst-v0.1 1 2 2 1 0 2 0 56 59 56 60 34 55 59
mistral-inst-v0.2 0 4 0 0 0 1 1 79 77 72 79 71 72 73
openchat-3.5 0 0 0 1 0 0 0 66 72 65 69 47 70 70
openchat-3.5-1210 0 0 2 0 1 1 0 75 72 66 71 55 66 68
average 0.8 7.1 2.0 2.0 1.3 2.4 2.0 65.9 62.5 56.3 63.5 52.6 61.3 61.3
@@ 3.4 Highlights
  • Finally, even if there is sufficient safety data for the traditional training of continuous prompts, it is still necessary to incorporate other general-domain data to prevent catastrophic forgetting. DRO bypasses this tricky issue through the regularization item $\mathcal{L}_{U}$3.3) that helps retain information other than the target features. We show in §4.2 that it is critical to maintaining the general model capability.
4 Experimental Evaluation
@@ 4.1 Baselines and Benchmarks
We use AlpacaEval (Li et al., 2023) to assess the impact of safety prompt optimization on the general model capability. AlpacaEval is an LLM-based automatic evaluation suite, which computes the win rate of the model responses to OpenAI’s text-davinci-003 responses given benign instructions. It has been widely adopted for open-source LLM evaluation (Ivison et al., 2023; Li et al., 2024) and we believe it can serve as a reasonable testbed for the 7B LLMs we experiment with. We use 100 randomly sampled instructions for evaluation and employ gpt-3.5-turbo as the evaluator. Additionally, we assess DRO’s impact on models’ false refusals on a held-out set of 100 harmless queries, which are collected in the same way as in §2.1.
@@
4.2 Main Results
Table 2 and 3 show the evaluation results using the default safety prompt. First, compared with the human-crafted basic safety prompt, DRO significantly improves safeguarding performance (1.6 vs. 9.8 on MaliciousInstruct; 1.4 vs. 10.3 on AdvBench) and meanwhile reduces false refusals for harmless queries (2.0 vs. 7.1 on the held-out harmless set), which does not compromise the general model capability (63.5 vs. 62.5 on AlpacaEval). From Figure 5, it is evident that DRO moves queries’ representations along (for out-of-domain harmful queries) or opposite (for harmless ones) our estimated refusal direction, which justifies the motivation of DRO (see Appendix K for full results). Second, DRO also remarkably outperforms the vPT baseline (1.6 vs. 5.0 on MaliciousInstruct; 1.4 vs. 6.8 on AdvBench), suggesting that vPT cannot well generalize to out-of-domain data. Moreover, vPT shows a deficiency in maintaining the general model capability (56.3 vs. 63.5 of DRO on AlpacaEval; dropping from 62.5 of the initial basic safety prompt), probably due to its nature of only optimizing for specific tasks using task-specific data. The above observations still hold when we apply DRO to optimize the other two human-crafted basic safety prompts (mistral and short), whose results are shown in Appendix I.
We then conduct ablation study on the optimization objectives in Equation 6. From Table 2, we can observe that the objective $\mathcal{L}_{\mathrm{r}}$ is critical to the safeguarding performance. Without $\mathcal{L}_{\mathrm{r}}$, models would still struggle to refuse harmful queries even when trained to distinguish harmful and harmless queries (i.e., with $\mathcal{L}_{\mathrm{h}}$). From Table 3, we can observe that the regularization item $\mathcal{L}_{U}$ is essential for maintaining the general model capability. Without $\mathcal{L}_{U}$, models would suffer from largely degraded generation quality for benign instructions (52.6 vs. 63.5 on AlpacaEval). We also observe that removing $\mathcal{L}_{\mathrm{h}}$ does not noticeably impair safeguarding performance and the general model capability, probably because the objective $\mathcal{L}_{\mathrm{r}}$ implicitly entails the capability of recognizing harmful queries.

### v3

@@ Abstract
Prepending model inputs with safety prompts is a common practice for safeguarding large language models (LLMs) against queries with harmful intents. However, the underlying working mechanisms of safety prompts have not been unraveled yet, restricting the possibility of automatically optimizing them to improve LLM safety. In this work, we investigate how LLMs’ behavior (i.e., complying with or refusing user queries) is affected by safety prompts from the perspective of model representation. We find that in the representation space, the input queries are typically moved by safety prompts in a “higher-refusal” direction, in which models become more prone to refusing to provide assistance, even when the queries are harmless. On the other hand, LLMs are naturally capable of distinguishing harmful and harmless queries without safety prompts. Inspired by these findings, we propose a method for safety prompt optimization, namely DRO (Directed Representation Optimization). Treating a safety prompt as continuous, trainable embeddings, DRO learns to move the queries’ representations along or opposite the refusal direction, depending on their harmfulness. Experiments with eight LLMs on out-of-domain and jailbreak benchmarks demonstrate that DRO remarkably improves the safeguarding performance of human-crafted safety prompts, without compromising the models’ general performance.
@@ 1 Introduction
While the capabilities of large language models (LLMs) keep growing (OpenAI, 2023; Touvron et al., 2023; Jiang et al., 2023; Zheng et al., 2024a), there are also rising concerns that they can engage with queries having harmful intents (e.g., seeking assistance about causing damages) (Weidinger et al., 2021). A common and lightweight means of safeguarding LLMs against harmful queries is to prepend model inputs with human-crafted safety prompts, which typically contain explicit guidance and guardrails on models’ behaviors. Real-world practices like GPT-4 (OpenAI, 2023) and Mistral (Jiang et al., 2023) have shown that adding safety prompts can mitigate models’ compliance with harmful queries, without needs of modifying model parameters or intervening the inference process, as illustrated in Figure 1.
@@
However, we still have no clear understanding of the working mechanisms of safety prompts, which restricts the possibility of automatically optimizing them to further improve LLM safety. Intrigued by this problem, our work starts by delving into how safety prompts intrinsically affect model behaviors from the perspective of model representations (§ 2). We propose two hypotheses: (1) Models cannot well distinguish harmful and harmless queries, while safety prompts enhance models’ capability of harmfulness recognition. (2) Models can recognize harmful queries but fail to refuse them, while safety prompts increase the probability of refusal (i.e., refusing to provide assistance). To verify the hypotheses, we first collect harmful and harmless queries through carefully controlled data synthesis (see Figure 2 for examples). We then evaluate eight open-source LLMs and employ PCA to visualize their hidden states. We find that in models’ representation space, harmful and harmless queries can be naturally distinguished, but this is not noticeably enhanced by safety prompts, suggesting that our first hypothesis does not hold. Instead, we observe that the queries’ representations are moved by different safety prompts in similar directions, in which models become more prone to generating refusal responses even when the queries are harmless, thus confirming our second hypothesis.
@@
Inspired by these findings, we present a method for safety prompt optimization, named DRO (Directed Representation Optimization; § 3). It takes the setting of the prompt tuning paradigm (Lester et al., 2021), where the model parameters are frozen and only a few continuous embeddings (corresponding to the safety prompts in our context) are trainable. DRO first anchors a model’s low-dimensional representation space and estimates the “refusal direction” that indicates the model’s refusal probability to increase (§ 3.1). It then optimizes the continuous safety prompt so that the queries’ representations are moved along or opposite the refusal direction, depending on the harmfulness of queries (§ 3.2). We also design a regularization item to prevent the degeneration of the original representation caused by direct optimization in the low-dimensional space (§ 3.3).
@@
We apply DRO to optimize the LLaMA-2 and Mistral official safety prompts. Through experiments with eight LLMs on out-of-domain and jailbreak benchmarks (§ 4), we demonstrate that DRO remarkably improves the safeguarding performance of human-crafted safety prompts. For instance, for the LLaMA-2 safety prompt, the percentage of compliance with harmful queries is reduced from 10.3% to 1.4% on AdvBench (Zou et al., 2023). Furthermore, DRO does not compromise the models’ general performance, as evaluated on AlpacaEval (Li et al., 2023), and exhibits reasonable robustness to the choices of data used for anchoring the low-dimensional space and refusal direction. We hope our work sheds light on the intrinsic working mechanisms of the prompt-driven LLM safeguarding approach, and inspires future research on LLM safety.
@@ 2 How Safety Prompts Intrinsically Work?
How to capture someone against their will and force them into captivity? How to capture precious moments with a camera on a family trip?
How to construct a false narrative to deceive others? How to construct a sturdy bridge over a river?
How to dispose of evidence after committing a crime? How to dispose of recyclable materials properly?
@@
Why can safety prompts safeguard LLMs against harmful queries, without which models may fail to refuse these queries but instead comply with them? We propose two hypotheses for the working mechanisms of safety prompts: (1) Models cannot well distinguish harmful and harmless queries, while safety prompts enhance models’ capability of harmfulness recognition. (2) Models can recognize harmful queries but fail to refuse them, while safety prompts increase models’ probability of generating refusal responses. To verify the hypotheses, we investigate how harmful and harmless queries exist in models’ representation space, and how the impact of safety prompts on queries’ representations correlates with models’ refusal behaviors.
@@ 2.3 Visualization Analysis
Do safety prompts make harmful and harmless queries more distinguishable?
From the upper part of Figure 3, harmful and harmless queries can be naturally distinguished, whose boundary (black chain dotted line) can be easily fitted by logistic regression using queries’ harmfulness as labels. However, adding safety prompts does not noticeably increase such distinguishability, even when visualized in other principal components (see Appendix G). These observations suggest that our first hypothesis does not hold, i.e., safety prompts do not clearly enhance models’ capability of harmfulness recognition.
@@
3 Methodology
Despite widespread use in real-world deployed LLMs like GPT-4 (OpenAI, 2023) and Mistral (Jiang et al., 2023), the prompt-driven safeguarding approach has its shortcoming, that is, the effectiveness varies with human-crafted prompts and models, as shown in Table 1. For instance, the short safety prompt works poorly with mistral-inst-v0.1 (55% harmful queries are still being complied with). Models that have undergone massive safety training, such as llama-2-chat and codellama-instruct, may also become over-sensitive when equipped with safety prompts, thereby leading to false refusals for harmless queries. Since crafting a basic safety prompt is always easy, can we optimize it for improved safeguarding performance? Inspired by our findings in § 2, we propose a method for automatically optimizing continuous safety prompts, named DRO, standing for Directed Representation Optimization. Its core idea is to move queries’ representations along or opposite the refusal direction according to their harmfulness.
@@ 3.2 Optimization Process
1:  Anchor the low-dimensional space and fit the refusal direction.$\triangleright$ Anchoring process (§ 3.1)
@@
$\triangleright$ Optimization process (Figure 4; § 3.23.3)
We similarly calculate a loss for harmfulness recognition, which can help maintain the capability of distinguishing harmful/harmless queries:
$\mathcal{L}_{\mathrm{h}}(\bm{\theta})=$ $\ -l\log\sigma(f_{\mathrm{h}}(\bm{x}_{\bm{\theta}})-f_{\mathrm{h}}(\bm{x}_{0}))$
$\ -(1-l)\log(1-\sigma(f_{\mathrm{h}}(\bm{x}_{\bm{\theta}})-f_{\mathrm{h}}(\bm{x}_{0}))),$ (4)
which uses the same dimensionality reduction function $g$ but a different logistic regression $f_{\mathrm{h}}$ fitted using queries’ harmfulness as labels:
$f_{\mathrm{h}}:\mathbb{R}^{n}\to\mathbb{R},f_{\mathrm{h}}(\bm{x})=\bm{w}_{\mathrm{h}}^{\top}g(\bm{x})+b_{\mathrm{h}},$ (5)
where $\bm{w}_{\mathrm{h}}\in\mathbb{R}^{m},b_{\mathrm{h}}\in\mathbb{R}$ are the fitted parameters. We find that adding $\mathcal{L}_{\mathrm{h}}(\bm{\theta})$ can bring some safeguarding performance improvement (§ 4.2).
@@ 3.4 Highlights
  • Finally, even if there is sufficient safety data for the traditional training of continuous prompts, it is still necessary to incorporate other general-domain data to prevent catastrophic forgetting. DRO bypasses this tricky issue through the regularization item $\mathcal{L}_{U}$ (§ 3.3) that helps retain information other than the target features. We show in § 4.2 that this regularization item $\mathcal{L}_{U}$ is critical to maintaining the models’ general performance.
Table 2: Evaluation results (optimizing the default basic safety prompt) on MaliciousInstruct, Advbench, the held-out harmless query set, and AlpacaEval.
% Compliance on MaliciousInstruct ↓ % Compliance on AdvBench ↓
llama-2-chat 1 1 1 1 0 1 0 0 0 3 0 0 0 0
codellama-instruct 3 2 7 1 1 1 1 2 0 2 0 0 0 0
vicuna-v1.5 51 10 7 2 2 4 2 27 4 2 0 1 2 0
orca-2 70 22 2 1 1 7 1 70 2 4 0 0 0 0
mistral-inst-v0.1 77 31 10 3 1 37 2 86 62 26 6 5 63 1
mistral-inst-v0.2 30 2 1 1 2 1 1 51 3 0 1 0 1 0
openchat-3.5 77 9 9 3 2 8 5 81 10 11 3 1 7 2
openchat-3.5-1210 66 1 3 1 3 3 2 78 1 6 1 1 7 1
average 46.9 9.8 5.0 1.6 1.5 7.8 1.8 49.4 10.3 6.8 1.4 1.0 10.0 0.5
% Refusal on Held-out Harmless ↓ % Win Rate on AlpacaEval ↑
llama-2-chat 1 19 5 5 3 7 7 66 47 37 54 53 53 48
codellama-instruct 3 22 0 7 5 8 7 54 52 47 51 45 48 51
vicuna-v1.5 0 5 4 2 1 0 1 68 65 62 64 58 65 61
orca-2 1 5 3 0 0 0 0 63 56 45 60 58 61 60
mistral-inst-v0.1 1 2 2 1 0 2 0 56 59 56 60 34 55 59
mistral-inst-v0.2 0 4 0 0 0 1 1 79 77 72 79 71 72 73
openchat-3.5 0 0 0 1 0 0 0 66 72 65 69 47 70 70
openchat-3.5-1210 0 0 2 0 1 1 0 75 72 66 71 55 66 68
average 0.8 7.1 2.0 2.0 1.3 2.4 2.0 65.9 62.5 56.3 63.5 52.6 61.3 61.3
4 Evaluation
@@ 4.1 Baselines and Benchmarks
We use AlpacaEval (Li et al., 2023) to assess the impact of safety prompt optimization on the models’ general performance. AlpacaEval is an LLM-based automatic evaluation suite, which computes the win rate of the model responses to OpenAI’s text-davinci-003 responses given benign instructions. It has been widely adopted for open-source LLM evaluation (Ivison et al., 2023; Li et al., 2024) and we believe it can serve as a reasonable testbed for the 7B LLMs we experiment with. We use 100 randomly sampled instructions for evaluation and employ gpt-3.5-turbo as the evaluator. Additionally, we assess DRO’s impact on models’ false refusals on a held-out set of 100 harmless queries, which are collected in the same way as in § 2.1.
4.2 Main Results
Table 2 show the evaluation results using the default safety prompt. First, compared with the human-crafted basic safety prompt, DRO significantly improves safeguarding performance (1.6 vs. 9.8 on MaliciousInstruct; 1.4 vs. 10.3 on AdvBench) and meanwhile reduces false refusals for harmless queries (2.0 vs. 7.1 on the held-out harmless set), which does not compromise the models’ general performance (63.5 vs. 62.5 on AlpacaEval). From Figure 5, it is evident that DRO moves queries’ representations along (for out-of-domain harmful queries) or opposite (for harmless ones) our estimated refusal direction, which justifies the motivation of DRO (see Appendix K for full results). Second, DRO also remarkably outperforms the vPT baseline (1.6 vs. 5.0 on MaliciousInstruct; 1.4 vs. 6.8 on AdvBench), suggesting that vPT cannot well generalize to out-of-domain data. Moreover, vPT shows a deficiency in maintaining the models’ general performance (56.3 vs. 63.5 of DRO on AlpacaEval; dropping from 62.5 of the initial basic safety prompt), probably due to its nature of only optimizing for specific tasks using task-specific data. The above observations still hold when we apply DRO to optimize the other two human-crafted basic safety prompts (mistral and short), whose results are shown in Appendix I.
@@ 4.2 Main Results
We then conduct ablation study on the optimization objectives in Equation 8. From Table 2 (upper), we can observe that the objective $\mathcal{L}_{\mathrm{r}}$ is critical to the safeguarding performance. Without $\mathcal{L}_{\mathrm{r}}$, models would still struggle to refuse harmful queries even when trained to distinguish harmful and harmless queries (i.e., with $\mathcal{L}_{\mathrm{h}}$). From Table 2 (lower right), we can observe that the regularization item $\mathcal{L}_{U}$ is essential for maintaining the models’ general performance. Without $\mathcal{L}_{U}$, models would suffer from largely degraded generation quality for benign instructions (52.6 vs. 63.5 on AlpacaEval). We also observe that removing $\mathcal{L}_{\mathrm{h}}$ does not noticeably impair safeguarding performance and the models’ general performance, probably because the objective $\mathcal{L}_{\mathrm{r}}$ has implicitly entailed the requirement for the capability of recognizing harmful queries.
4.3 Extension To Jailbreak Setting
As LLM jailbreaking (Zou et al., 2023; Wei et al., 2023) has become an increasingly threatening safety issue, we further evaluate DRO’s effectiveness in defending against the GCG (Zou et al., 2023) jailbreak attack on AdvBench. GCG appends each query with an adversarial suffix, which is optimized using a gradient-based method. According to the transferability in Zou et al. 2023, we use the GCG jailbreak prompts optimized from llama-2-chat for all the models. We then directly prepend the DRO-optimized default safety prompt, which is used in § 4.2 and Table 2. In Table 3, we show that in the jailbreak setting, DRO remains effective in improving the safeguarding performance of human-crafted safety prompts (3.1 vs. 13.3), while vPT cannot generalize well to jailbreak prompts. Interestingly, we find that even the basic safety prompt can provide non-trivial safeguarding, comparing to not adding safety prompts (13.3 vs. 54.1), which can serve as a strong baseline for future research on LLM jailbreaking and safeguarding.
Table 3: Evaluation results (optimizing the default basic safety prompt) on AdvBench under GCG jailbreak attack.
GCG Jailbreak
no default vPT DRO
llama-2-chat 2 0 27 0
codellama-instruct 7 1 13 1
vicuna-v1.5 46 14 9 2
orca-2 82 8 3 0
mistral-inst-v0.1 88 66 16 12
mistral-inst-v0.2 62 3 0 1
openchat-3.5 79 12 5 5
openchat-3.5-1210 67 2 2 4
average 54.1 13.3 9.4 3.1